OpenStudy (anonymous):

determine if lim as x approaches 0 (x^6+8x^2)/(4x^7+9x^11) exists and if it does find its value. Would it be positive or negative infinity?

OpenStudy (anonymous):

what do you get when you replace \(x\) bu \(0\)?

OpenStudy (anonymous):

\[\frac{x^6+8x^2}{4x^7+9x^11}=\frac{x^4+8}{4x^5+9x^9}\] \]

OpenStudy (anonymous):

limit does not exist because numerator does not go to zero, whilst the denominator does